# SEIR models with host heterogeneity: theoretical aspects and applications to seasonal influenza dynamics

### v1

@@ 6.3 Results
One may also observe that the estimated values of $\mathcal{R}_{0}$ are very similar between the two models (though consistently larger for the model with heterogeneity) and quite larger than the corresponding estimates by [Trentini2022] for the same years.
@@
Figure 4 illustrates the fitted curves for selected seasons. We highlight the 2016–17 and 2023–24 seasons because these are the years for which the $R^{2}$ values of the two models differ the most. In contrast, 2019–20 stands out as the only season where the homogeneous model achieves a higher $R^{2}$ than the heterogeneous one.”
@@ 7 Discussion
The paper focuses on incorporating host heterogeneity in susceptibility and infectiousness into SEIR epidemiological models. The method used is based on the results by [DiekmannInaba2023], which have been adapted specifically for the SEIR model; we believe that the presentation for this specific setting may be easier to follow than the general treatment in [DiekmannInaba2023].
Following this path, in the case of a generic distribution of susceptibility, one obtains system (8) when susceptibility and infectiousness are independent, whereas system (LABEL:sys_bar2b) when they are correlated. The cases of a Gamma (already considered by [DiekmannInaba2023]) and of a Beta distribution of susceptibility are then illustrated.
@@
The result is consistent with the findings by [Novozhilov2008], which followed a very different approach and did not consider the case of correlated susceptibility and infectiousness.
In the case where susceptibility is Gamma-distributed, we also show that the final size decreases as a function of the variance of the Gamma distribution, both with independent and correlated infectiousness. Furthermore, the final size is smaller with correlated than with independent infectiousness for the same variance of the distribution. In biological terms, greater variability in the host population dampens epidemic spread, as the more resistant individuals limit chains of transmission.
Empirical fits to Italian influenza surveillance data indicate that the model with Gamma-distributed heterogeneity in susceptibility is sufficiently flexible to adapt to the dynamics of all normal seasons of influenza infections. Indeed, the homogeneous SEIR model can also accurately describe the dynamics of seasonal influenza, but only assuming that a considerable fraction of individuals were already immune to the circulating strain at season start; however, estimates, based on serological data, of the fraction of the population immune to the circulating strains at season start [Lunelli2013] are much lower than the estimates for $R(0)$ in Table 1(a). Thus, in order to fit the available data, the homogeneous SEIR model must make unrealistic assumptions about pre-existing immunity.
We believe, rather, that epidemic outbreak models should incorporate not only known sources of heterogeneity, such as age, spatial location, or social class, but also unobserved variation in susceptibility, represented by an additional parameter (such as the $p$ used here). Such intrinsic heterogeneity may reflect not only biological differences among individuals but also differences in behaviour that affect the probability of getting infected, as in [BrittonBallTrapman2020]. Heterogeneity may also be tied to the evolution of strains, so that individuals may have been last infected by strains with different similarities to the currently circulating strain [Andreasen2003a, Britton2025].
@@
It is easy to extend the current model to allow for multiple groups connected through a given contact matrix. From a practical point of view, the question is whether one can assume the same distribution of susceptibility for each group or whether it would be necessary to allow for different distributions, leading to a multiplication of parameters. We plan to explore this issue by extending the present analysis to include age structure [Trentini2022] in the model and in the available data, as well as known vaccination rates and effectiveness [Calabro2025].
@@ Appendix A Proofs
Let $G_{k}(w)=\Phi_{k-1}(w)-\gamma/\beta\log(w)$ for $k=1$ or 2. Clearly $\lim\limits_{w\to 0^{+}}G(w)=+\infty$, while $G(1)=1$, Furthermore
@@
$\Phi_{1}(\bar{s}_{\infty})-\frac{\Phi_{2}(1)}{\mathcal{R}_{0}}\ln(\bar{s}_{\infty})=1.$
@@
$H(w):=\frac{\Phi_{1}(w)}{\Phi_{2}(1)}-\frac{\ln(w)}{R_{0}}-\frac{1}{\Phi_{2}(1)}$
@@
$S^{\prime}_{\infty}(p)=\dfrac{R_{0}\cdot\left(S_{\infty}(p)-1\right)-\left(S_{\infty}(p)\right)^{-\tfrac{1}{p}}\cdot\ln\left(S_{\infty}(p)\right)}{p^{2}G_{x}(S_{\infty}(p),p)}.$ (28)
@@
$R_{0}\cdot\left(1-S_{\infty}(p)\right)<-\ln\left(S_{\infty}(p)\right)\cdot\left(S_{\infty}(p)\right)^{-\tfrac{1}{p}}.$
@@
$p\cdot\left(\left(S_{\infty}(p)\right)^{-\tfrac{1}{p}}-1\right)<-\ln\left(S_{\infty}(p)\right)\cdot\left(S_{\infty}(p)\right)^{-\tfrac{1}{p}},$
@@
$\left(p+\ln\left(S_{\infty}(p)\right)\right)\cdot\left(S_{\infty}(p)\right)^{-\tfrac{1}{p}}<p\\ \iff 1+\frac{1}{p}\ln\left(S_{\infty}(p)\right)<\left(S_{\infty}(p)\right)^{\tfrac{1}{p}}\\ \iff 1+\ln\left(S_{\infty}(p)\right)^{\frac{1}{p}}<\left(S_{\infty}(p)\right)^{\tfrac{1}{p}}.$
@@
${S^{*}_{\infty}}^{\prime}(p)=\dfrac{p^{2}\cdot{S^{*}_{\infty}}^{\tfrac{1}{p}}-p^{2}-(p+1)\ln{S^{*}_{\infty}}+R_{0}\cdot{S^{*}_{\infty}}^{1+\tfrac{2}{p}}\cdot\ln{S^{*}_{\infty}}}{p^{2}\mathcal{R}_{0}{S^{*}_{\infty}}^{\tfrac{1}{p}}F_{x}(S^{*}_{\infty}(p),p)}.$ (30)
@@
$p^{2}\cdot\left({S^{*}_{\infty}}^{\tfrac{1}{p}}-1\right)>(p+1)\ln{S^{*}_{\infty}}-R_{0}\cdot{S^{*}_{\infty}}^{1+\tfrac{2}{p}}\cdot\ln{S^{*}_{\infty}}.$ (31)
@@
$\left(1-{S^{*}_{\infty}}^{\tfrac{1}{p}}\right)<-\frac{p+1}{p^{2}}\cdot\ln\left({S^{*}_{\infty}}\right)\cdot\Bigg(1-\frac{\left({S^{*}_{\infty}}^{-\tfrac{1}{p}}-1\right)\cdot{S^{*}_{\infty}}^{1+\tfrac{2}{p}}}{1-{S^{*}_{\infty}}^{1+\tfrac{1}{p}}}\Bigg).$
@@
$1-{S^{*}_{\infty}}^{\tfrac{1}{p}}<-\ln\left({S^{*}_{\infty}}^{\tfrac{1}{p}}\right)=-\dfrac{1}{p}\ln\left({S^{*}_{\infty}}\right),$
@@
$\ln\left({S^{*}_{\infty}}\right)>-\frac{p+1}{p}\cdot\ln\left({S^{*}_{\infty}}\right)\cdot\Bigg(1-\frac{\left({S^{*}_{\infty}}^{-\tfrac{1}{p}}-1\right)\cdot{S^{*}_{\infty}}^{1+\tfrac{2}{p}}}{1-{S^{*}_{\infty}}^{1+\tfrac{1}{p}}}\Bigg).\\$
Since $\ln\left({S^{*}_{\infty}}\right)$ is negative, after multiplying by $p/(p+1)$, this is equivalent to

### v2

@@ 6.3 Results
One may also observe that the estimated values of $\mathcal{R}_{0}$ are very similar between the two models (though consistently larger for the model with heterogeneity) and quite larger than the corresponding estimates by [23] for the same years.
@@
Figure 7 illustrates the fitted curves for selected seasons. We highlight the 2016–17 and 2023–24 seasons because these are the years for which the $R^{2}$ values of the two models differ the most. In contrast, 2019–20 stands out as the only season where the homogeneous model achieves a higher $R^{2}$ than the heterogeneous one.
@@ 7 Discussion
The paper focuses on incorporating host heterogeneity in susceptibility and infectiousness into SEIR epidemiological models. The method used is based on the results by [11], which have been adapted specifically for the SEIR model; we believe that the presentation for this specific setting may be easier to follow than the general treatment in [11].
Following this path, in the case of a generic distribution of susceptibility, one obtains system (8) when susceptibility and infectiousness are independent, whereas system (LABEL:sys_bar2b) when they are correlated. The cases of a gamma distribution (already considered by [11]) and of a Beta distribution of susceptibility are then illustrated.
@@
The result is consistent with the findings by [18], which followed a very different approach and did not consider the case of correlated susceptibility and infectiousness.
In the case where susceptibility is gamma-distributed, we also show that the final size decreases as a function of the variance of the gamma distribution, both with independent and correlated infectiousness. Furthermore, the final size is smaller with correlated than with independent infectiousness for the same variance of the distribution. In biological terms, greater variability in the host population dampens epidemic spread, as the more resistant individuals limit chains of transmission. Notice that assuming that infectiousness is proportional to susceptibility may cover the case in which the difference in susceptibility reflects differences in behaviour rather than in immunological response to infection challenges. In such a case, in fact, individuals with more contacts, will be, on the one hand, more likely to get infected, and, on the other hand, more infectious, assuming that differences in behaviour are retained also after infection.
Empirical fits to Italian influenza surveillance data indicate that the model with gamma-distributed heterogeneity in susceptibility is sufficiently flexible to adapt to the dynamics of all normal seasons of influenza infections. Indeed, the homogeneous SEIR model can also accurately describe the dynamics of seasonal influenza, but only assuming that a considerable fraction of individuals were already immune to the circulating strain at season start; however, estimates, based on serological data, of the fraction of the population immune to the circulating strains at season start [16] are much lower than the estimates for $R(0)$ in Table 1(a). Thus, in order to fit the available data, the homogeneous SEIR model must make unrealistic assumptions about pre-existing immunity.
Notice that we only considered retrospective fits to the infection data over a whole season. Predicting the time course of an influenza season is notoriously an extremely different task to which many different methods have been applied, with mixed success [24, 19].
We believe, rather, that epidemic outbreak models should incorporate not only known sources of heterogeneity, such as age, spatial location, or social class, but also unobserved variation in susceptibility, represented by an additional parameter (such as the $p$ used here). Such intrinsic heterogeneity may reflect not only biological differences among individuals but also differences in behaviour that affect the probability of getting infected, as in [7]. Heterogeneity may also be tied to the evolution of strains, so that individuals may have been last infected by strains with different similarities to the currently circulating strain [4, 8].
@@
It is easy to extend the current model to allow for multiple groups connected through a given contact matrix. From a practical point of view, the question is whether one can assume the same distribution of susceptibility for each group or whether it would be necessary to allow for different distributions, leading to a multiplication of parameters. We plan to explore this issue by extending the present analysis to include age structure [23] in the model and in the available data, as well as known vaccination rates and effectiveness [9].
@@ Appendix A Proofs
Let $G_{k}(w)=\Phi_{k-1}(w)-\gamma/\beta\log(w)$ for $k=1$ or 2. Clearly $\lim\limits_{w\to 0^{+}}G_{k}(w)=+\infty$, while $G_{k}(1)=1$ for $k=1$ or 2. Furthermore
@@
$\Phi_{1}(\bar{s}_{\infty})-\frac{\Phi_{2}(1)}{\mathcal{R}_{0}}\log(\bar{s}_{\infty})=1.$
@@
$H(w):=\frac{\Phi_{1}(w)}{\Phi_{2}(1)}-\frac{\log(w)}{R_{0}}-\frac{1}{\Phi_{2}(1)}$
@@
$S^{\prime}_{\infty}(p)=\dfrac{R_{0}\cdot\left(S_{\infty}(p)-1\right)-\left(S_{\infty}(p)\right)^{-\tfrac{1}{p}}\cdot\log\left(S_{\infty}(p)\right)}{p^{2}G_{x}(S_{\infty}(p),p)}.$ (29)
@@
$R_{0}\cdot\left(1-S_{\infty}(p)\right)<-\log\left(S_{\infty}(p)\right)\cdot\left(S_{\infty}(p)\right)^{-\tfrac{1}{p}}.$
@@
$p\cdot\left(\left(S_{\infty}(p)\right)^{-\tfrac{1}{p}}-1\right)<-\log\left(S_{\infty}(p)\right)\cdot\left(S_{\infty}(p)\right)^{-\tfrac{1}{p}},$
@@
$\left(p+\log\left(S_{\infty}(p)\right)\right)\cdot\left(S_{\infty}(p)\right)^{-\tfrac{1}{p}}<p\\ \iff 1+\frac{1}{p}\log\left(S_{\infty}(p)\right)<\left(S_{\infty}(p)\right)^{\tfrac{1}{p}}\\ \iff 1+\log\left(S_{\infty}(p)\right)^{\frac{1}{p}}<\left(S_{\infty}(p)\right)^{\tfrac{1}{p}}.$
@@
${S^{*}_{\infty}}^{\prime}(p)=\dfrac{p^{2}\cdot{S^{*}_{\infty}}^{\tfrac{1}{p}}-p^{2}-(p+1)\log{S^{*}_{\infty}}+R_{0}\cdot{S^{*}_{\infty}}^{1+\tfrac{2}{p}}\cdot\log{S^{*}_{\infty}}}{p^{2}\mathcal{R}_{0}{S^{*}_{\infty}}^{\tfrac{1}{p}}F_{x}(S^{*}_{\infty}(p),p)}.$ (31)
@@
$p^{2}\cdot\left({S^{*}_{\infty}}^{\tfrac{1}{p}}-1\right)>(p+1)\log{S^{*}_{\infty}}-R_{0}\cdot{S^{*}_{\infty}}^{1+\tfrac{2}{p}}\cdot\log{S^{*}_{\infty}}.$ (32)
@@
$\left(1-{S^{*}_{\infty}}^{\tfrac{1}{p}}\right)<-\frac{p+1}{p^{2}}\cdot\log\left({S^{*}_{\infty}}\right)\cdot\Bigg(1-\frac{\left({S^{*}_{\infty}}^{-\tfrac{1}{p}}-1\right)\cdot{S^{*}_{\infty}}^{1+\tfrac{2}{p}}}{1-{S^{*}_{\infty}}^{1+\tfrac{1}{p}}}\Bigg).$
@@
$1-{S^{*}_{\infty}}^{\tfrac{1}{p}}<-\log\left({S^{*}_{\infty}}^{\tfrac{1}{p}}\right)=-\dfrac{1}{p}\log\left({S^{*}_{\infty}}\right),$
@@
$\log\left({S^{*}_{\infty}}\right)>-\frac{p+1}{p}\cdot\log\left({S^{*}_{\infty}}\right)\cdot\Bigg(1-\frac{\left({S^{*}_{\infty}}^{-\tfrac{1}{p}}-1\right)\cdot{S^{*}_{\infty}}^{1+\tfrac{2}{p}}}{1-{S^{*}_{\infty}}^{1+\tfrac{1}{p}}}\Bigg).\\$
Since $\log\left({S^{*}_{\infty}}\right)$ is negative, after multiplying by $p/(p+1)$, this is equivalent to
@@
Proof of Proposition 5.4.
Remembering $p=1/V$, the thesis is equivalent to
$S_{\infty}(p)<\frac{1+p\bar{x}}{p+1}.$ (33)
To prove (33), first of all notice that, manipulating (23), one sees that $S_{\infty}(p)$ can be seen as the solution in $(0,1)$ of
$G(x):=\log(x)+p\log\left(1+\frac{\mathcal{R}_{0}}{p}(1-x)\right)=0.$ (34)
From the properties of $G(\cdot)$, it follows that (33) is equivalent, using $R_{0}(1-\bar{x})=-\log(\bar{x})$, to
$\log\left(\dfrac{1+p\bar{x}}{p+1}\right)+p\log\left(1-\frac{\log\bar{x}}{p+1}\right)>0.$ (35)
We will prove that (35) holds for any $0<\bar{x}<1.$ Let
$g(x):=\log(1+px)-\log(p+1)+p\log\left(1-\frac{\log x}{p+1}\right).$
It is clear that $g(1)=0.$ To prove that $g(x)>0$ for $x\in(0,1)$, we will check that $g(x)$ is decreasing in $\left(0,1\right)$. Differentiating yields
$g^{\prime}(x)=p\left(\frac{1}{1+px}-\frac{1}{x(1+p)\left(1-\dfrac{\log x}{p+1}\right)}\right)$
$=p\left(\frac{1}{1+px}-\frac{1}{x+px-x\log x}\right).$
Considering $p>0$, we check that
$\frac{1}{1+px}<\frac{1}{x+px-x\log x}\iff 1+px>x+px-x\log x$
$\iff x\log x>x-1,$
which holds in $(0,1).$
Appendix B Supplementary figures
Refer to caption Refer to caption
Figure 8: Left panel: Attack ratio at the end of an epidemic for system (7), plotted as a function of the variance with $R_{0}=3$; curves are shown when susceptibility $X$ follows the gamma distribution, or the Beta distributions with either the parameter choices in (19) or (20) and when $X$ is either equal to 0 or to $V+1$ (bimodal distribution). Right panel: Same as left panel but for system (8)–(LABEL:sys_bar2b).
Refer to caption Refer to caption
Figure 9: Left panel: Attack ratio at the end of an epidemic for systems (7) (8)–(LABEL:sys_bar2b), plotted as a function of the variance with $R_{0}=3$; curves are shown when susceptibility $X$ follows the gamma distribution, or the Beta distributions with the parameter choices in (19). Right panel: Attack ratio at the end of an epidemic for systems (7) (8)–(LABEL:sys_bar2b), plotted as a function of the variance with $R_{0}=3$; curves are shown when susceptibility $X$ follows the gamma distribution, or the Beta distributions with the parameter choices in (20).